OpenStudy (anonymous):

Complex Numbers (-10 + 15i) divide by 5i

OpenStudy (jdoe0001):

\(\bf -10+15i \div 5i\implies \cfrac{-10+15i}{5i}\implies \cfrac{-10}{5i}+\cfrac{15i}{5i}\)

OpenStudy (anonymous):

\[\frac{ -10+15\iota }{5\iota }*\frac{ \iota }{\iota }=\frac{ -10\iota+15 \iota ^{2} }{5\iota ^{2} }\] \[=\frac{ 10\iota-15 }{ -5 }=-2\iota+3=3-2\iota \]